# AVOCADO: Adaptive Optimal Collision Avoidance driven by Opinion

### v1

@@ Abstract
We present AVOCADO (AdaptiVe Optimal Collision Avoidance Driven by Opinion), a novel navigation approach to address holonomic robot collision avoidance when the degree of cooperation of the other agents in the environment is unknown. AVOCADO departs from a Velocity Obstacle’s formulation akin to the Optimal Reciprocal Collision Avoidance method. However, instead of assuming reciprocity, AVOCADO poses an adaptive control problem that aims at adapting in real-time to the cooperation degree of other robots and agents. Adaptation is achieved through a novel nonlinear opinion dynamics design that relies solely on sensor observations. As a by-product, based on the nonlinear opinion dynamics, we propose a novel method to avoid the deadlocks under geometrical symmetries among robots and agents. Extensive numerical simulations show that AVOCADO surpasses existing geometrical, learning and planning-based approaches in mixed cooperative/non-cooperative navigation environments in terms of success rate, time to goal and computational time. In addition, we conduct multiple real experiments that verify that AVOCADO is able to avoid collisions in environments crowded with other robots and humans.
@@ I Introduction
Robot navigation in dynamic environments is the task of moving a robot from one location to another while avoiding collisions with obstacles that are in motion around the environment[1]. Collision avoidance is particularly challenging when the dynamic obstacles are agents (humans, other robots, etc.) with their own intent and degree of cooperation, because it is hard for robots to predict their future motion and effectively plan to avoid potential collisions[2, 3]. The ability of avoiding collisions in crowded dynamic environments is a low-level feature that is of key importance in robotic applications such as social robotics[4, 5], robots in the wild [6, 7, 8] or autonomous driving[9, 10].
Existing solutions (Section II) make one or more of the following assumptions: (i) the degree of cooperation between the robot and an agent is known a priori (e.g., the robot knows that the agent is polite and will leave space for the robot to cross), restricting the solution to predetermined cooperative settings[11, 12, 13, 14]; (ii) the robot can communicate with other robots and/or agents[15, 16, 17, 18], which might not be possible during deployment due to a lack of cooperation, communication delays, packet losses or disturbances from external factors; (iii) the computational capabilities of the robot are sufficient to do inference on learning-based policies or long-term horizon planning[19, 20, 21, 22], which is hard to ensure in low-cost onboard platforms if we take into account that collision avoidance is just a low-level feature among a myriad of desired high-level and computationally demanding capabilities (e.g., multi-robot active perception or semantic mapping).
@@
To cope with all these issues, the main contribution of this work is AVOCADO (AdaptiVe Optimal Collision Avoidance Driven by Opinion), a novel method for collision avoidance in cooperative and non-cooperative dynamic environments (Section IV). AVOCADO is built upon a geometrical formulation based on the classical Velocity Obstacle (VO, [23, 24]) framework (Section III). However, different from other VO-based approaches that assume either no cooperation or reciprocity [25, 11, 26], we consider that the degree of cooperation between the robot and the agent is unknown and do not assume any predetermined behavior for the agent or communication infrastructure. To address uncertainty on the degree of cooperation, we propose a novel nonlinear opinion dynamics adaptive scheme that enables the robot to adapt to the agent’s behavior in real-time from just its onboard sensor observations. Furthermore, we exploit the nonlinear opinion dynamics adaptive law to develop a novel method that overcomes the geometrical deadlocks coming from symmetries that suffer existing VO-based methods. We validate our proposal with simulations (Section V) and real experiments (Section VI), showing that AVOCADO surpasses exiting collision avoidance approaches in environments that mix cooperative and non-cooperative agents in terms of avoided collisions while featuring an inexpensive computational cost. The scenarios include examples that combine real robots and humans in tight spaces, concluding (Section VII) that AVOCADO is a promising alternative for low-cost robot navigation in crowded environments where agents exhibit unknown degrees of cooperation.
The code will be made available upon acceptance.
@@ II Related work
Research on collision avoidance was primarily dominated by geometrical approaches. Departing from robot observations, methods like potential fields[27], dynamic window[28], inevitable collision states[29] or social forces[30] compute control actions that steer the robot towards collision-free regions; they are simple to compute but lack formal performance guarantees, leading to over-conservative policies. By contrast, VO-based methods[31] preserve the computational simplicity but guarantee optimality in terms of deviation with respect to the desired predefined velocity either for holonomic or non-holonomic robots [26, 32, 33]. Specially relevant is ORCA[11], which builds a half-space geometrical constraint per agent that is incorporated in a linear program. ORCA and other methods based on Reciprocal VO (RVO) assume reciprocity [25], i.e, the robot and agent cooperate equally to avoid the collision, which is not the case in mixed cooperative/non-cooperative environments such the ones considered in this paper.
@@
To address uncertainty on the degree of cooperation of the agents, some solutions [34] remove the reciprocity assumption in ORCA and formulate an additional optimization problem. Instead of using optimization, ORCA can also be implemented as a low-level module integrated in a high-level learned policy that accounts for social and traffic cues[35]. Both approaches require 1-hop communication, which is not feasible in a non-cooperative environment due to the lack of cooperation among agents. Instead, AVOCADO is fully decentralized and only relies on the perception of the robot.
@@
The latter solution[35, 36] is a deep learning approach for collision avoidance in robot navigation[19, 20]. Either from a supervised[37, 38, 39, 40] or reinforcement learning perspective[41, 42, 43, 44, 45, 46], deep learning develop end-to-end policies that map the perception observations into control commands[13] or sub-goals that are fed into a low-level horizon planning controller[47]. For instance, long-short term memory networks can be combined with barrier certificates[48] to improve collision avoidance success rate and scalability. Nonetheless, deep learning approaches have a high memory and computational cost, often lack collision avoidance guarantees and fail in out-of-distribution settings.
@@
The lack of guarantees of purely learning-based methods is addressed by reachability methods[15], control barrier functions[49] and model predictive control[50]. Reachability-based approaches depart from the Hamilton-Jacobi equation[51] to build a safe set in the state space that fulfills all the desired constrains. Forward computation of the Hamilton-Jacobi equation is highly demanding and over-conservative[52], so learning-based alternatives[53, 54, 55] propose approximations that trade computation and performance guarantees. On the other hand, control barrier functions formulate an optimization problem that finds the closest control action to the desired one that satisfies a set of desired forward invariance constraints, stemming from a control stability definition of the desired safe set. It is not clear how to extend control barrier functions to dynamic obstacles with unknown behaviors[56] or communication-denied settings[57, 58]. Finally, model predictive control shares with reachability methods the use of a time-horizon propagation and with control barrier functions the optimization-based formulation[59]. Model predictive control is flexible because it enables to satisfy additional goals or constraints such as motion uncertainties[60] or local maps[21]; however, it has a significant computational burden that is only alleviated by reducing the time-horizon and the number of considered nearby agents, resulting in undesired local minima or oscillations. Different from these alternatives, AVOCADO is inexpensive to compute, is not over-conservative as it minimizes the deviation with respect to the desired velocity, and does not suffer from symmetry deadlocks and oscillations.
@@
One key characteristic of the aforementioned methods is that, in non-cooperative environments with unknown intents, a prediction model of the behavior of the other agents is required to effectively avoid collisions[61, 62]. The first option is to use learned predictors that directly obtain future trajectory states of the agents given current and past samples[63, 64]. The second option is to rely on a learned model[65], such that it can be integrated in a model predictive control program[22] to simultaneously optimize over the trajectories of the robot and the other agents. In both cases, the success of the collision avoidance module is subject to the accuracy of the complex prediction modules; instead, AVOCADO adapts to unknown degrees of cooperation through a novel nonlinear opinion dynamics adaptive law that is inexpensive to compute, does not require learning and is only based on perception observations.
@@
Opinion dynamics [66, 67, 68] originate from studies on how to model social interactions. The ineffectiveness of linear models to represent behaviors like saturation of information leads to nonlinear opinion dynamics[69]. Regarding collision avoidance, an important property of nonlinear opinion dynamics is that they ensure the existence of a bifurcation in the state-space, i.e., the existence of two simultaneous stable equilibrium points. This feature has been exploited [70] to design a collision avoidance method that avoids deadlocks, where the opinion represents the preference of the robot to move left or right; nevertheless, this approach does not guarantee collision avoidance. AVOCADO exploits nonlinear opinion dynamics to guarantee that the robot decides if the agent is cooperative or non-cooperative and in what degree, leaving the collision avoidance guarantees to a VO-based program.
@@ III Problem formulation
To adapt to the unknown degree of cooperation, we rely on the concept of nonlinear opinion dynamics[69]. Opinion dynamics emerge as a means of modeling interactive behaviors where nodes in a network can choose among a discrete set of options regarding one or more topics. Let $x_{j}\in[-1,1]$ be the opinion that node $j$ has about a certain topic. Typically, $x_{j}=0$ represents neutrality whereas $x_{j}=\{-1,1\}$ represent extreme opinions over a topic with two possible opinions. By interacting with other nodes, node $j$ can evolve its opinion. For a topic with two opinions, one possible nonlinear opinion dynamics design is
$\dot{x}_{j}=-d_{j}x_{j}+g(x_{j})f\left(a_{j}x_{j}+\sum_{k=1,k\neq j}^{\mathsf{M}}c_{jk}x_{k}\right)+b_{j}.$ (5)
In the equation above, $d_{j}>0$ tunes how fast the current opinion vanishes with time; $b_{j}\in\mathbb{R}$ is a bias that models inherent priorities that the node $j$ has on the topic; $g(x_{j})$ is an attention term that evolves with time and weights the importance of the influence of other nodes; $a_{j}>0,c_{jk}>0$ are consensus-like weights that model the exchange of opinions among the $\mathsf{M}>0$ nodes of the network; and $f(\bullet)$ is a nonlinear function, such as $\mathsf{tanh}$ or $\mathsf{sigmoid}$, that bounds the influence of the interactions. More importantly, the nonlinear function in the opinion dynamics enforces a bifurcation phenomenon. The bifurcation phenomenon is characterized by the appearance of a pitchfork bifurcation in the evolution of the state variable $x_{j}$ that leads to two non-zero stable equilibrium, each of them associated to one of the opinions. Whether the state evolves to one or another depends on the interactions between nodes. For more details on this, we refer the reader to [69]. In relation with this work, the inherent bifurcation property of nonlinear opinion dynamics and the conceptual relationships between the opinion and the parameters of Eq. (5) motivate their use to estimate (“build an opinion”) on the unknown degree of cooperation of the other agents.
@@ IV-A Optimal Collision Avoidance for Unknown Degrees of Cooperation
We propose AVOCADO to address the robot collision avoidance problem in multi-agent dynamic environments. The first step towards deriving AVOCADO is to consider the geometry defined by $\mathsf{VO}_{i}^{\tau}$ for all $i\in\{1,\dots,\mathsf{N}\}$. Let assume that the relative velocity $\mathbf{v}_{r}^{\mathsf{pre}}-\mathbf{v}_{i}\in\mathsf{VO}_{i}^{\tau}$ for some agent $i$. This means that, if robot and agent $i$ follow $\mathbf{v}_{r}^{\mathsf{pre}}$ and $\mathbf{v}_{i}$ respectively, then collision will happen at any time in $[0,\tau]$. The velocity $\mathbf{v}_{r}^{\mathsf{pre}}$ is provided by some higher-level planner, while $\mathbf{v}_{i}$ is known from onboard sensor measurements. The latter can be measured or estimated in real-time from onboard cameras or LiDAR [73, 74]. Since cooperation of agent $i$ cannot be assumed, then the robot needs to choose a new velocity $\mathbf{v}_{r}^{*}$ to avoid the collision.
Let $\partial\mathsf{VO}_{i}^{\tau}$ denote the boundary of the velocity obstacle set $\mathsf{VO}_{i}^{\tau}$. Then, the minimum change in velocity that makes $\mathbf{v}_{r}^{\mathsf{pre}}-\mathbf{v}_{i}\notin\mathsf{VO}_{i}^{\tau}$ is given by the vector $\mathbf{u}_{i}$
@@
Given $\mathsf{OCA}_{i}$ for all $i\in\{1,\ldots,\mathsf{N}\}$, AVOCADO selects the closest velocity to $\mathbf{v}_{r}^{\mathsf{pre}}$ that respects all the constraints sets. Formally, this is given by the following linear program
@@
We remark that (8) is a natural extension of ORCA to settings where the degree of cooperation of the agents is not necessarily reciprocal. The challenge, in this sense, is how to estimate $\alpha_{i}$ for all perceived agents $i\in\mathcal{N}$ from local sensing. Next, we describe how AVOCADO exploits opinion dynamics to adapt online to the degree of cooperation of each agent.

### v4

@@ Abstract
We present AVOCADO (AdaptiVe Optimal Collision Avoidance Driven by Opinion), a novel navigation approach to address holonomic robot collision avoidance when the robot does not know how cooperative the other agents in the environment are. AVOCADO departs from a Velocity Obstacle’s (VO) formulation akin to the Optimal Reciprocal Collision Avoidance method. However, instead of assuming reciprocity, it poses an adaptive control problem to adapt to the cooperation level of other robots and agents in real time. This is achieved through a novel nonlinear opinion dynamics design that relies solely on sensor observations. As a by-product, we leverage tools from the opinion dynamics formulation to naturally avoid the deadlocks in geometrically symmetric scenarios that typically suffer VO-based planners. Extensive numerical simulations show that AVOCADO surpasses existing motion planners in mixed cooperative/non-cooperative navigation environments in terms of success rate, time to goal and computational time. In addition, we conduct multiple real experiments that verify that AVOCADO is able to avoid collisions in environments crowded with other robots and humans.
@@
©This paper has been accepted for publication at IEEE Transactions on Robotics. Please, when citing the paper, refer to the official manuscript with the following DOI: 10.1109/TRO.2025.3552350.
@@ I Introduction
Robot navigation in dynamic environments is the task of moving a robot from one location to another while avoiding collisions with obstacles that are in motion around the environment[1]. Collision avoidance is particularly challenging when the dynamic obstacles are agents (humans, other robots, etc.), and the problem of accounting for the effect of the robot motion on the other agents is still unsolved[2]. Agents have their own intent and they exhibit different degrees of cooperation, i.e. they adjust their behavior to avoid collision with the robot to a certain extent that is unknown to the robot. Thus, it is hard for robots to predict their future motion and effectively plan to avoid potential collisions[3, 4]. The ability of avoiding collisions in crowded dynamic environments is a low-level feature that is of key importance in robotic applications such as social robotics[5, 6], robots in the wild [7, 8, 9] or autonomous driving[10, 11].
Existing solutions (Section II) make one or more of the following assumptions: (i) the degree of cooperation between the robot and an agent is known a priori (e.g., the robot knows that the agent is going to collaborate in the collision avoidance and how), restricting the solution to predetermined cooperative settings[12, 13, 14, 15]; (ii) the robot can communicate with other robots and/or agents[16, 17, 18, 19], which might not be possible during deployment due to a lack of cooperation, communication delays, packet losses or disturbances from external factors; (iii) the computational capabilities of the robot are sufficient to do inference on learning-based policies or long-term horizon planning[20, 21, 22, 23], which is hard to ensure in low-cost onboard platforms if we take into account that collision avoidance is just a low-level feature among a myriad of desired high-level and computationally demanding capabilities (e.g., multi-robot active perception or semantic mapping [24, 25]).
@@
To cope with all these issues, the main contribution of this work is AVOCADO (AdaptiVe Optimal Collision Avoidance Driven by Opinion), a novel method for collision avoidance in cooperative and non-cooperative dynamic environments (Section IV). AVOCADO is built upon a geometrical formulation based on the classical Velocity Obstacle (VO, [26, 27]) framework (Section III). However, different from other VO-based approaches that assume either no cooperation or reciprocity [28, 12, 29], we consider that the degree of cooperation between the robot and the agent is unknown and do not assume any predetermined behavior for the agent or communication infrastructure. To address uncertainty on the degree of cooperation, we propose a novel nonlinear opinion dynamics adaptive scheme that enables the robot to adapt to the agent’s behavior in real-time from just its onboard sensor observations. Furthermore, we exploit the nonlinear opinion dynamics adaptive law to develop a method that avoids the geometrical deadlocks coming from symmetries that suffer some existing VO-based methods. We validate our proposal with simulations (Section V) and real experiments (Section VI), showing that AVOCADO surpasses exiting collision avoidance approaches in environments that mix cooperative and non-cooperative agents in terms of avoided collisions while featuring an inexpensive computational cost. The scenarios include examples that combine real robots and humans in tight spaces, concluding (Section VII) that AVOCADO is a promising alternative for low-cost robot navigation in crowded environments where agents exhibit unknown degrees of cooperation.
The code of the work is implemented in C++/Python and can be accessed in the repository of the work [^1] https://github.com/dmartinezbaselga/AVOCADO, as well as the videos of the experiments.
@@ II Related work
Research on collision avoidance was primarily dominated by geometrical approaches. Departing from robot observations, methods originally only considered static obstacles, as the Dynamic Window Approach [30]. Others like potential fields[31], inevitable collision states [32] or social forces[33] compute control actions that reactively lead the robot towards collision-free regions. They are simple to compute but lack formal performance guarantees, leading to over-conservative policies. By contrast, VO-based methods[34] preserve the computational simplicity but guarantee optimality in terms of deviation with respect to the desired predefined velocity either for holonomic or non-holonomic robots [29, 35, 36]. Reciprocal VO (RVO) extends VO by assuming reciprocity in the collision avoidance [28], i.e, it assumes that agents encountered in the environment cooperate equally to avoid the collision. Specially relevant is Optimal Reciprocal Collision Avoidance ORCA [12], which builds a half-space geometrical constraint per agent that is incorporated in a linear program. Nevertheless, the reciprocal avoidance assumption may not be applied in mixed cooperative/non-cooperative environments, which are the ones considered in this paper.
@@
Some solutions [37] remove the reciprocity assumption in ORCA and formulate an additional optimization problem to reduce the time the agents take to reach the goal. Instead of using optimization, ORCA can also be implemented as a low-level module integrated in a high-level learned policy that accounts for social and traffic cues[38]. Both approaches require 1-hop communication, which is not feasible in a non-cooperative environment due to the lack of cooperation among agents. Instead, AVOCADO is fully decentralized and only relies on the perception of the robot.
@@
Many recent methods use deep learning for robot navigation. Either from a supervised[20, 39, 40, 41, 42] or reinforcement learning perspective[43, 44, 21, 45, 46, 47, 48], deep learning develop end-to-end policies that map the perception observations into control commands[14] or sub-goals that are fed into a low-level horizon planning controller[49]. For instance, Long-Short Term Memory (LSTM) networks can be combined with barrier certificates[50] to improve collision avoidance success rate and scalability. There are also works that combine ORCA scenario knowledge with reinforcement learning [38, 51]. Nonetheless, deep learning approaches have a high memory and computational cost, often lack collision avoidance guarantees and fail in out-of-distribution settings.
@@
The lack of guarantees of purely learning-based methods is addressed by reachability methods[16], control barrier functions[52] and Model Predictive Control (MPC)[53]. Reachability-based approaches depart from the Hamilton-Jacobi equation[54] to build a safe set in the state space that fulfills all the desired constrains. Forward computation of the Hamilton-Jacobi equation is highly demanding and over-conservative[55], so learning-based alternatives[56, 57, 58] propose approximations that trade computation and performance guarantees. On the other hand, control barrier functions formulate an optimization problem that finds the closest control action to the desired one that satisfies a set of desired forward invariance constraints, stemming from a control stability definition of the desired safe set. It is not clear how to extend control barrier functions to dynamic obstacles with unknown behaviors[59] or communication-denied settings[60, 61]. Finally, MPC shares with reachability methods the use of a time-horizon propagation and with control barrier functions the optimization-based formulation[62]. MPC is flexible because it enables to satisfy additional goals or constraints such as motion uncertainties[63] or local maps[22]. However, it has a significant computational burden that is only alleviated by reducing the time-horizon and the number of considered nearby agents, resulting in undesired local minima or oscillations. Different from these alternatives, AVOCADO is inexpensive to compute, is not over-conservative as it minimizes the deviation with respect to the desired velocity, and does not suffer from symmetry deadlocks and oscillations.
@@
One key characteristic of the aforementioned methods is that, in non-cooperative environments with unknown intents, a prediction model of the behavior of the other agents is required to effectively avoid collisions[64, 65]. The first option is to use learned predictors that directly obtain future trajectory states of the agents given current and past samples[66, 67]. The second option is to rely on a learned model[68], such that it can be integrated in a MPC program[23] to simultaneously optimize over the trajectories of the robot and the other agents. In both cases, the success of the collision avoidance module is subject to the accuracy of the complex prediction modules; instead, AVOCADO adapts to unknown degrees of cooperation through a novel nonlinear opinion dynamics adaptive law that is inexpensive to compute, does not require learning and is only based on current perception observations, without relying on prediction of future trajectories.
@@
Opinion dynamics [69, 70, 71] originate from studies on how to model social interactions. The ineffectiveness of linear models to represent behaviors like saturation of information leads to nonlinear opinion dynamics[72, 73]. They have been applied in a wide variety of fields, such as explaining political polarization [74] or perception and reaction to epidemics [75]; lately, they have been applied in robotic problems, mainly combining them with game theory for cooperative and non-cooperative multi-agent decision making [76, 77, 78] or multi-robot task allocation [79]. Regarding collision avoidance, there is an important property of nonlinear opinion dynamics. They ensure the existence of a bifurcation in the state-space, i.e., the existence of two simultaneous stable equilibrium points. This feature has been exploited [80] to design a collision avoidance method that avoids deadlocks, where the opinion represents the preference of the robot to move left or right. A recent work [81] extends opinion dynamics from the planar to the circle space to represent, e.g., the heading angle of a robot. Nevertheless, these approaches do not guarantee collision avoidance nor consider other motion patterns than only changing the heading angle. AVOCADO exploits nonlinear opinion dynamics to guarantee that the robot decides if the agent is cooperative or non-cooperative and in what degree, leaving the collision avoidance guarantees to a VO-based program.
@@ III Problem formulation
To adapt to the unknown degree of cooperation, we rely on the concept of nonlinear opinion dynamics[72]. Opinion dynamics emerge as a means of modeling interactive behaviors where nodes in a network can choose among a discrete set of options regarding one or more topics. Let $o_{j}\in[-1,1]$ be the opinion that node $j$ has about a certain topic. Typically, $o_{j}=0$ represents neutrality whereas $o_{j}=\{-1,1\}$ represent extreme opinions over a topic with two possible opinions. By interacting with other nodes, node $j$ can evolve its opinion. For a topic with two opinions, one possible nonlinear opinion dynamics design is
$\dot{o}_{j}=-d_{j}o_{j}+g(o_{j})f\left(a_{j}o_{j}+\sum_{k=1,k\neq j}^{\mathsf{M}}c_{jk}o_{k}\right)+b_{j}.$ (5)
In the equation above, $d_{j}>0$ tunes how fast the current opinion vanishes with time; $b_{j}\in\mathbb{R}$ is a bias that models inherent priorities that the node $j$ has on the topic; $g(o_{j})$ is an attention term that evolves with time and weights the importance of the influence of other nodes; $a_{j}>0,c_{jk}>0$ are consensus-like weights that model the exchange of opinions among the $\mathsf{M}>0$ nodes of the network; and $f(\bullet)$ is a nonlinear function, such as $\mathsf{tanh}$ or $\mathsf{sigmoid}$, that bounds the influence of the interactions. More importantly, the nonlinear function in the opinion dynamics enforces a bifurcation phenomenon. The bifurcation phenomenon is characterized by the appearance of a pitchfork bifurcation in the evolution of the state variable $o_{j}$ that leads to two non-zero stable equilibrium, each of them associated to one of the opinions. Whether the state evolves to one or another depends on the interactions between nodes. For more details on this, we refer the reader to [72]. In relation with this work, the inherent bifurcation property of nonlinear opinion dynamics and the conceptual relationships between the opinion and the parameters of Eq. (5) motivate their use to estimate (“build an opinion”) on the unknown degree of cooperation of the other agents.
@@ IV-A Optimal Collision Avoidance for Unknown Degrees of Cooperation
We propose AVOCADO to address the robot collision avoidance problem in multi-agent dynamic environments. The first step towards deriving AVOCADO is to consider the geometry defined by $\mathsf{VO}_{i}$ for all $i\in\{1,\dots,\mathsf{N}\}$. Let assume that the relative velocity $\mathbf{v}_{r}^{\mathsf{pre}}-\mathbf{v}_{i}\in\mathsf{VO}_{i}$ for some agent $i$. This means that, if robot and agent $i$ follow $\mathbf{v}_{r}^{\mathsf{pre}}$ and $\mathbf{v}_{i}$ respectively, then collision will happen at any time in $[0,\tau]$. The velocity $\mathbf{v}_{r}^{\mathsf{pre}}$ is provided by some higher-level planner, while $\mathbf{v}_{i}$ is known from onboard sensor measurements. The latter can be measured or estimated in real-time from onboard cameras or LiDAR [84, 85]. Since cooperation of agent $i$ cannot be assumed, then the robot needs to choose a new velocity $\mathbf{v}_{r}^{*}$ to avoid the collision.
Let $\partial\mathsf{VO}_{i}$ denote the boundary of the velocity obstacle set $\mathsf{VO}_{i}$. Then, the minimum change in velocity that makes $\mathbf{v}_{r}^{\mathsf{pre}}-\mathbf{v}_{i}\notin\mathsf{VO}_{i}$ is given by the vector $\mathbf{u}_{i}$
@@
Given $\mathsf{OCA}_{i}$ for all $i\in\{1,\ldots,\mathsf{N}\}$, AVOCADO selects the closest velocity to $\mathbf{v}_{r}^{\mathsf{pre}}$ that respects all the constraints sets at the same time. Formally, this is given by the following linear program
@@
Problem (8) is guaranteed to have a solution unless the intersection of the $\mathsf{OCA}_{i}$ sets is empty. In this case, the robot computes the velocity that minimizes the distance with the half-planes $\mathsf{OCA}_{i}\quad\forall i\in\mathcal{N}$, i.e., it chooses the “least unsafe” velocity in the absence of a velocity that guarantees collision avoidance in that instant (see Section 5.3 of [12] for further details). We remark that (8) is a natural extension of ORCA to settings where the degree of cooperation of the agents is not necessarily reciprocal. Fig. 3 shows a representation of the constrained set (8b).
Figure 3: Representation of the intersection set specified by the $\mathsf{OCA}_{i}$ half-planes of $4$ agents for different values of $\alpha_{i}$. Note that $\alpha_{i}$ may be different for each of the agents, but it is considered the same in this figure for a simple visualization. If the intersection set of all half-planes is not empty, the problem is guaranteed to have a solution.
The challenge is how to estimate $\alpha_{i}$ for all perceived agents $i\in\mathcal{N}$ from local sensing. Next, we describe how AVOCADO exploits opinion dynamics to adapt online to the degree of cooperation of each agent.